OpenStudy (anonymous):

Express the complex number in trigonometric form. -2i

OpenStudy (jdoe0001):

\(\large { -2i\implies 0-2i\implies \begin{array}{cccllll} (0&,&-2)\\ x&&y \end{array}\qquad r=\sqrt{x^2+y^2}\qquad (r,\theta) }\)

OpenStudy (dumbcow):

\[z = a+bi = r (\cos \theta +i \sin \theta)\] \[a^2 + b^2 = r^2\] \[\frac{b}{a} = \tan \theta\] for this problem: a = 0 b = -2

OpenStudy (jdoe0001):

now if you graph ( 0, -2) where would you end up as an angle \(\theta?\)

OpenStudy (jdoe0001):



OpenStudy (anonymous):

Is it 270 degrees?

OpenStudy (jdoe0001):

 yeap